# Flamelet Connection to Turbulence Kinetic Energy Dissipation Rate

### v1

@@ 3 Basic Relations for Rotational Flamelet
Consider the transformation from the non-rotating coordinate system with space coordinates $x,y,z$ and corresponding velocity components $u,v,w$ to the system rotating about the $z$-axis through the origin with the coordinates $\xi,\chi,z$ and velocity components $u_{\xi},u_{\chi},z$. The transformation displayed in Figure 1 is made from the Newtonian frame with rotating material (due to vorticity) to a rotating, non-Newtonian frame. Let the vorticity direction be the $z$ direction in an orthogonal framework. Any $x,y$ plane contains the directions of scalar gradients, major principal axis for compressive strain, and a principal axis for tensile strain. Note that the $x,y,z$ directions are not correlated with coordinates on the resolved scale. $\omega$ is the vorticity magnitude on this sub-grid scale. $x,y,z$ are transformed to $\xi,\chi,z$ wherein the material rotation is removed from the $\xi,\chi$ plane by having it rotate at angular velocity $d\theta/dt=\omega/2$ relative to $x,y$. Here, $\theta$ is the angle between the $x$ and $\xi$ axes and simultaneously the angle between the $y$ and $\chi$ axes. We take the sub-grid domain to be sufficiently small to consider a uniform value of $\omega$ across it.
@@
$\frac{\partial u_{\chi}}{\partial\chi}$ $=$ $\frac{\partial v}{\partial\chi}cos\theta-\frac{\partial u}{\partial\chi}cos\theta\;;\;\frac{\partial u_{\chi}}{\partial\xi}=0=\frac{\partial v}{\partial\xi}cos\theta-\frac{\partial u}{\partial\xi}sin\theta-\frac{\omega}{2}$ (6)
From Equations (3) and (4), multiplications by $cos\theta$ or $sin\theta$, addition of equations, and use of $cos^{2}\theta+sin^{2}\theta=1$ yields
$\frac{\partial u_{\xi}}{\partial\xi}cos\theta=\frac{\partial u}{\partial\xi}+\frac{\omega sin\theta}{2}\;\;;\;\;\frac{\partial u_{\chi}}{\partial\chi}cos\theta=\frac{\partial v}{\partial\chi}+\frac{\omega sin\theta}{2}$ (7)
Solution of four linear algebraic equations given by Equations (3) and (4) yields
$\frac{\partial u}{\partial\xi}$ $=$ $\frac{\partial u}{\partial x}cos\theta+\frac{\partial u}{\partial y}sin\theta\;\;;\;\;\frac{\partial u}{\partial\chi}=\frac{\partial u}{\partial y}cos\theta-\frac{\partial u}{\partial x}sin\theta\;\;;$
$\frac{\partial v}{\partial\xi}$ $=$ $\frac{\partial v}{\partial x}cos\theta+\frac{\partial v}{\partial y}sin\theta\;\;;\;\;\frac{\partial v}{\partial\chi}=\frac{\partial v}{\partial y}cos\theta-\frac{\partial v}{\partial x}sin\theta$ (8)
Use of Equations (8) to substitute into Equation (2) followed by division with $cos\theta$ yields
$\frac{\partial u_{\xi}}{\partial\xi}$ $=$ $\frac{\partial u}{\partial x}+\Big{(}\frac{\partial u}{\partial y}+\frac{\omega}{2}\Big{)}tan\theta$
@@
The inflow should be independent of $\theta$ position. That occurs if $\partial v/\partial x=-\partial u/\partial y=\omega/2$. This implies that the shear strain rate $S_{xy}=[\partial u/\partial y+\partial v/\partial x]/2=0$. Since $w$ depends only on $z$ and $u$ and $v$ are independent of $z$ for the inflow, $S_{xz}=0,S_{yz}=0$ as well. Thereby, only normal strain rate occurs in the inflow for either coordinate system.
With the aim towards establishing scaling relations, the flamelet-related material in the following several sections is applied to incoming counterflow at some distance from the reaction zone. There, density, other scalar properties, strain rates can be considered as uniform; velocity gradients and scalar gradients will differ from values at the filtered scales of the resolved flow. Our scaling from resolved-scale values for RANS or LES will impose boundary conditions on the flamelet but will not prescribe directly the values in or near the reaction zone; in this critical manner, the approach differs from the Flamelet Progress Variable method.
4 Divergence and Curl of Velocity
Given the relation for the vorticity in the original reference frame,
$\frac{\partial v}{\partial x}-\frac{\partial u}{\partial y}=\omega$ (10)
and the relations in Equation (9), it follows that the divergence of velocity remains the same in both frames. Namely,
@@ 4 Divergence and Curl of Velocity
Identical dilatation will be felt in both coordinate frames.
For flow into the flamelet counterflow, $u_{\xi}$ is only a function of $\xi$ and $u_{\chi}$ is only dependent on $\chi$, and the velocity derivatives $\partial u_{\xi}/\partial\chi$ and $\partial u_{\chi}/\partial\xi$ are each zero. Thereby, the curl of the velocity for the inflow in the rotating frame is zero.
$\frac{\partial u_{\chi}}{\partial\xi}-\frac{\partial u_{\xi}}{\partial\chi}=0$ (12)
That is, the vorticity does not appear explicitly for the inflow in the description within the rotating frame. It can have the important implicit effect through the centrifugal acceleration. In the flamelet interior core with varying density, velocity derivatives indicating shear strain and vorticity can develop because $u_{\xi}(\xi,\chi)$ and $w(\chi,z)$ can appear. However, since $u_{\xi}$ is antisymmetric in $\xi$ and $w$ is antisymmetric in $z$, the circulations for the regions with total vorticity in the $\xi$- or $\chi$-directions or added vorticity in the $z$-direction will each be zero.
@@ 5 Viscous Dissipation and Enstrophy
$+\boldmath{\Big{[}\Big{(}\frac{\partial u}{\partial z}\Big{)}^{2}+\Big{(}\frac{\partial v}{\partial z}\Big{)}^{2}+\Big{(}\frac{\partial w}{\partial x}\Big{)}^{2}+\Big{(}\frac{\partial w}{\partial y}\Big{)}^{2}+2\frac{\partial u}{\partial z}\frac{\partial w}{\partial x}+2\frac{\partial v}{\partial z}\frac{\partial w}{\partial y}\Big{]}}$
@@
Consider now that $\omega/2=\partial v/\partial x=-\partial u/\partial y$. Then, Equation (9) yields $\partial u/\partial x=\partial u_{\xi}/\partial\xi$ and $\partial v/\partial y=\partial u_{\chi}/\partial\chi$. Also, the last three terms on the first line of Equation (17) become $\omega^{2}/4+\omega^{2}/4-2\omega^{2}/4=0$. Now, for the inflow, the viscous dissipation can be described by
@@
If a portion or all of the flow is divergence free, the effect of bulk viscosity is absent and Equations (15,16,5) may be written as
@@
$\frac{1}{\rho}\frac{\partial^{2}p}{\partial x_{i}^{2}}$ $=$ $-\frac{\partial}{\partial x_{i}}\Big{(}u_{j}\frac{\partial u_{i}}{\partial x_{j}}\Big{)}+\mu\frac{\partial^{2}}{\partial x_{j}^{2}}\Big{(}\frac{\partial u_{i}}{\partial x_{i}}\Big{)}$ (20)
$=$ $-\frac{\partial u_{j}}{\partial x_{i}}\frac{\partial u_{i}}{\partial x_{j}}-u_{j}\frac{\partial}{\partial x_{j}}\frac{\partial u_{i}}{\partial x_{i}}+\mu\frac{\partial^{2}}{\partial x_{j}^{2}}\Big{(}\frac{\partial u_{i}}{\partial x_{i}}\Big{)}$
@@
Realize that, in the rotating frame the centrifugal acceleration is given as $\vec{a}=(\xi\omega^{2}/4,\chi\omega^{2}/4)$. Thereby, $\nabla\cdot\vec{a}=\omega^{2}/2$. Since viscous dissipation maintains the same value in both systems, the implication is that the Laplacian of the pressure is identical in both systems. In order to maintain a pressure maximum and thereby a stagnation point, the centrifugal effect must be overcome by the effect of the strain rate; i.e., $\omega^{2}<A_{ij}A_{ij}$ where $A_{ij}\equiv\partial u_{i}/\partial x_{j}$. Note that $\Phi/\mu$ is independent of the coefficient of viscosity. The balance here relates to strain rate and not viscous action. Actually, the viscous effect was removed by the divergence operation.
Here, we have a strong argument for the importance of the balance of vorticity magnitude and strain-rate magnitude or, more precisely, the balance between dissipation rate and enstrophy. There are many studies on these quantities, especially for incompressible flows (Johnson, 2021; Johnson & Wilczek, 2023; Yeung et al., 2012). If enstrophy $\omega^{2}$ becomes too large compared to the dissipation rate divided by dynamic viscosity $\Phi/\mu$, the pressure Laplacian becomes positive, thereby disallowing the flamelet counterflow possibility since the required stagnation point located at a pressure maxima cannot occur. This vital issue of balance cannot be addressed in a formulation that neglects vorticity and only addresses strain rate. For incompressible turbulent flow, this balance has been addressed by Yeung et al. (2012) who find that the temporal statistics for both quantities present similar magnitudes. They find large values of each parameter tend to occur together with $\Phi/\mu>\omega^{2}$ for 60-65 percent of the data. Although the study does not include dilatation, it shows support for the concept of mixing and reaction occurring in rotating counterflows provided by the turbulent eddies. In cases where combustion has mixing as rate-controlling, the higher-strain-rate (and thereby smaller) eddies would provide the highest burning rates (provided the mixing rate does not exceed the reaction rate).
@@
$\omega^{2}=\frac{\partial u_{i}}{\partial x_{j}}\frac{\partial u_{i}}{\partial x_{j}}-2\frac{\partial u_{i}}{\partial x_{j}}\frac{\partial u_{j}}{\partial x_{i}}+\frac{\partial u_{j}}{\partial x_{i}}\frac{\partial u_{j}}{\partial x_{i}}=2\Big{[}\frac{\partial u_{i}}{\partial x_{j}}\frac{\partial u_{i}}{\partial x_{j}}-\frac{\partial u_{i}}{\partial x_{j}}\frac{\partial u_{j}}{\partial x_{i}}\Big{]}$ (21)
Then, with account for bulk viscosity, Equations (16) and (21) yield
$\epsilon=\frac{\Phi}{\rho}=\nu\Big{(}2\frac{\partial u_{i}}{\partial x_{j}}\frac{\partial u_{i}}{\partial x_{j}}-\frac{2}{3}\Big{[}\frac{\partial u_{k}}{\partial x_{k}}\Big{]}^{2}-\frac{\omega^{2}}{2}\Big{)}$ (22)
@@
where the definition $S_{1}+S_{2}=1$ has been entered. The implication of Equation (5) is that if $\epsilon$ were known from the resolved scale, a relation between $S^{*}$ and $S_{1}$ is established for the rotational flamelet. This relation could be useful for the Flamelet Progress Variable model as well. For example, with $S_{1}=1/2$, we find $S^{*}=\sqrt{\epsilon/(3\nu)}$, its peak value. As $S_{1}$ varies in the range $0\leq S_{1}\leq 1$, $S^{*}$ varies narrowly between $\sqrt{\epsilon/(4\nu)}$ and $\sqrt{\epsilon/(3\nu)}$. In the range $-1\leq S_{1}\leq 0$ (with inward swirl), $S^{*}$ varies more strongly between $\sqrt{\epsilon/(12\nu)}$ and $\sqrt{\epsilon/(3\nu)}$.
Viscous dissipation actually occurs over a range of the smallest scales and not totally at the Kolmogorov scale. Therefore, we have reason to multiply $\epsilon$ in Equation (22) by the coefficient $C_{vd}$ where $C_{vd}<1$.
@@ 6 Kinetic Energy
With use of Equation (2), it can be readily found that
@@
For the incompressible counterflow structure, we can write
@@
The right side of Equation (23) requires knowledge of three quantities for rotational flamelet theory: $S^{*},S_{1},$ and vorticity $\omega$. The Kolmogorov scale can be used to estimate the lengths. The kinetic energy per unit mass here is based on Kolmogorov scale velocity fluctuations; so, the value will be less than the statistical value for turbulence kinetic energy $k$. If a relation between the two kinetic energies is provided, knowledge of $k$ provides a relationship between those three quantities. Thereby, Equations (22) and (23) have potential to be used with knowledge of $\epsilon$ from resolved-scale computation to determine the needed inflow parameters $S^{*}$ and $\omega$ for the rotational flamelet model.
@@
We will use the relations from Pope, Page 185 (Pope, 2000); the Kolmogorov velocity $u_{\kappa}\approx(\epsilon\nu)^{1/4}$ and the Kolmogorov length $\kappa\approx(\nu^{3}/\epsilon)^{1/4}$. We will average kinetic energy over a volume equal to $\kappa^{3}$. The last term in Equation (23) shows antisymmetry, taking positive values in two quadrants and negative values of identical magnitudes in the other two quadrants of any $\xi,\chi$ plane. Thereby, they have no consequence for our averaging. Equation (23) now yields after averaging the right side and relating the left side to $\epsilon$,
@@
$C_{ke}$ is a nondimensional coefficient of $O(1)$ that appears in the averaging process for the kinetic energy term on the left side of Equations (23). It accounts for both approximations made in averaging and adjusted decrease in kinetic energy at the Kolmogorov scale because dissipation has occurred over a range of the smallest scales.

### v2

@@ 3 Basic Relations for Rotational Flamelet
Consider the rotational flamelet transformation from the non-rotating coordinate system with space coordinates $x,y,z$ and corresponding velocity components $u,v,w$ to the system rotating about the $z$-axis through the origin with the coordinates $\xi,\chi,z$ and velocity components $u_{\xi},u_{\chi},w$ [Sirignano2022a]. Note that in this section, the symbol $\chi$ refers to the rotating-frame coordinate, not to the scalar dissipation rate as discussed in previous sections. The transformation displayed in Figure 1 is made from the Newtonian frame with rotating material (due to vorticity) to a rotating, non-Newtonian frame. The vorticity direction is the $z$ direction in an orthogonal framework. Any $x,y$ plane contains the directions of scalar gradients, major principal axis for compressive strain, and a principal axis for tensile strain. Note that the $x,y,z$ directions are not correlated with coordinates on the resolved scale. $\omega$ is the vorticity magnitude on this sub-grid scale. $x,y,z$ are transformed to $\xi,\chi,z$ wherein the material rotation is removed from the $\xi,\chi$ plane by having it rotate at angular velocity $d\theta/dt=\omega/2$ relative to $x,y$. Here, $\theta$ is the angle between the $x$ and $\xi$ axes and simultaneously the angle between the $y$ and $\chi$ axes. The sub-grid domain is sufficiently small to consider a uniform value of $\omega$ across it.
@@
$\frac{\partial u_{\chi}}{\partial\chi}$ $=$ $\frac{\partial v}{\partial\chi}cos\theta-\frac{\partial u}{\partial\chi}sin\theta\;;\;\frac{\partial u_{\chi}}{\partial\xi}=0=\frac{\partial v}{\partial\xi}cos\theta-\frac{\partial u}{\partial\xi}sin\theta-\frac{\omega}{2}$ (6)
Equations (5) and (6) now can yield
$\frac{\partial u_{\xi}}{\partial\xi}$ $=$ $\frac{\partial u}{\partial x}+\Big(\frac{\partial u}{\partial y}+\frac{\omega}{2}\Big)tan\theta$ (7)
@@
Now, we focus on the inflow on either side of the counterflow induced by a compressive normal strain rate imposed on the rotating mass of fluid. Asymptotically with distance from the center of rotation, the inflow has uniform density, although exothermic reaction in the center of the rotating fluid will cause variable density locally. The divergence of velocity in either coordinate system will be zero for the incoming flow. The inflow should be independent of $\theta$ position. That occurs if $\partial v/\partial x=-\partial u/\partial y=\omega/2$. This implies that the shear strain rate $S_{xy}=[\partial u/\partial y+\partial v/\partial x]/2=0$. Since $w$ depends only on $z$ and $u$ and $v$ are independent of $z$ for the inflow, $S_{xz}=0$, $S_{yz}=0$ as well. Thereby, only normal strain rate occurs in the inflow for either coordinate system.
With the aim towards establishing scaling relations, the flamelet-related material is applied to incoming counterflow at some distance from the reaction zone. There, density, other scalar properties, and strain rates are considered as uniform. Velocity gradients and scalar gradients will differ from values at the filtered scales of the resolved flow. Our scaling from resolved-scale values for RANS or LES will impose boundary conditions on the flamelet but will not prescribe directly the values in or near the reaction zone; in this critical manner, the approach differs from the Flamelet Progress Variable method [Pierce].
Given the relations for the vorticity in the original reference frame and the transformed frame, we have
$\frac{\partial v}{\partial x}-\frac{\partial u}{\partial y}=\omega\;\;;\;\;\frac{\partial u_{\chi}}{\partial\xi}-\frac{\partial u_{\xi}}{\partial\chi}=0$ (9)
The dilation remains identical in both frames; thus, it follows that the divergence of velocity is unchanged in value. Namely,
@@
Furthermore, the divergence will be zero for the incoming flow; although density variation will occur in the flamelet interior due to exothermic reaction.
For flow into the flamelet counterflow at a distance from the interface of the two incoming streams, $u_{\xi}$ is only a function of $\xi$ and $u_{\chi}$ is only dependent on $\chi$, and the velocity derivatives $\partial u_{\xi}/\partial\chi$ and $\partial u_{\chi}/\partial\xi$ are each zero. That is, the vorticity does not appear explicitly for the inflow in the description within the rotating frame. It can have the important implicit effect through the centrifugal acceleration. In the flamelet interior core with varying density, velocity derivatives indicating shear strain and vorticity can develop because $u_{\xi}(\xi,\chi)$ and $w(\chi,z)$ can appear. However, since $u_{\xi}$ is antisymmetric in $\xi$ and $w$ is antisymmetric in $z$, the circulations for the regions with total vorticity in the $\xi$- or $\chi$-directions or added vorticity in the $z$-direction will each be zero.
@@ 4 Viscous Dissipation and Enstrophy
$+\Big[\Big(\frac{\partial u}{\partial z}\Big)^{2}+\Big(\frac{\partial v}{\partial z}\Big)^{2}+\Big(\frac{\partial w}{\partial x}\Big)^{2}+\Big(\frac{\partial w}{\partial y}\Big)^{2}+2\frac{\partial u}{\partial z}\frac{\partial w}{\partial x}+2\frac{\partial v}{\partial z}\frac{\partial w}{\partial y}\Big]$
@@
With $\omega/2=\partial v/\partial x=-\partial u/\partial y$, as shown above, Equations (7) and (8) yield $\partial u/\partial x=\partial u_{\xi}/\partial\xi$ and $\partial v/\partial y=\partial u_{\chi}/\partial\chi$. Also, the last three terms on the first line of Equation (15) become $\omega^{2}/4+\omega^{2}/4-2\omega^{2}/4=0$. Now, for the inflow, the viscous dissipation can be described by
@@
We consider the incoming streams to the flamelet which are divergence free and Equations (13,14, 16) may be written as
@@
$\frac{1}{\rho}\frac{\partial^{2}p}{\partial x_{i}^{2}}$ $=$ $-\frac{\partial}{\partial x_{i}}\Big(u_{j}\frac{\partial u_{i}}{\partial x_{j}}\Big)+\frac{\mu}{\rho}\frac{\partial^{2}}{\partial x_{j}^{2}}\Big(\frac{\partial u_{i}}{\partial x_{i}}\Big)$ (18)
$=$ $-\frac{\partial u_{j}}{\partial x_{i}}\frac{\partial u_{i}}{\partial x_{j}}-u_{j}\frac{\partial}{\partial x_{j}}\frac{\partial u_{i}}{\partial x_{i}}+\frac{\mu}{\rho}\frac{\partial^{2}}{\partial x_{j}^{2}}\Big(\frac{\partial u_{i}}{\partial x_{i}}\Big)$
@@
Realize that, in the rotating frame, the centrifugal acceleration is given as $\vec{a}=(\xi\omega^{2}/4,\chi\omega^{2}/4)$. Thereby, $\nabla\cdot\vec{a}=\omega^{2}/2$. Since viscous dissipation maintains the same value in both systems, the implication is that the Laplacian of the pressure is identical in both systems. In order to maintain a pressure maximum and thereby a stagnation point, the centrifugal effect must be overcome by the effect of the strain rate; i.e., $\omega^{2}<A_{ij}A_{ij}$ where $A_{ij}\equiv\partial u_{i}/\partial x_{j}$. Note that $\Phi/\mu$ is independent of the coefficient of viscosity. The balance here relates to strain rate and not viscous action. Actually, the viscous effect was removed by the divergence operation. Still, we use use common terminology that refers to $A_{ij}A_{ij}$ as “dissipation rate”.
Here, we have a strong argument for the importance of the balance of vorticity magnitude and strain-rate magnitude or, more precisely, the balance between dissipation rate and enstrophy. There are many studies on these quantities, especially for incompressible flows [PJohnson, JohnsonWilczek, Yeungetal]. If enstrophy $\omega^{2}$ becomes too large compared to the dissipation rate divided by dynamic viscosity $\Phi/\mu$, the pressure Laplacian becomes positive, thereby disallowing the flamelet counterflow possibility since the required stagnation point located at a pressure maxima cannot occur. This vital issue of balance cannot be addressed in a formulation that neglects vorticity and only addresses strain rate. For incompressible turbulent flow, this balance has been addressed by [Yeungetal] who find that the temporal statistics for both quantities present similar magnitudes. They find large values of each parameter tend to occur together with $\Phi/\mu>\omega^{2}$ for 60-65 percent of the data. Although the study does not include dilatation, it shows support for the concept of mixing and reaction occurring in rotating counterflows provided by the turbulent eddies. In cases where combustion has mixing as rate-controlling, the higher-strain-rate (and thereby smaller) eddies would provide the highest burning rates (provided the mixing rate does not exceed the reaction rate).
@@
$\omega^{2}=\Big[\frac{\partial u_{i}}{\partial x_{j}}\frac{\partial u_{i}}{\partial x_{j}}-\frac{\partial u_{i}}{\partial x_{j}}\frac{\partial u_{j}}{\partial x_{i}}\Big]$ (19)
Then, Equations (14) and (19) yield
$\epsilon=\frac{\Phi}{\rho}=\nu\Big(2\frac{\partial u_{i}}{\partial x_{j}}\frac{\partial u_{i}}{\partial x_{j}}-\frac{2}{3}\Big[\frac{\partial u_{k}}{\partial x_{k}}\Big]^{2}-\omega^{2}\Big)=\nu\left(2\frac{\partial u_{i}}{\partial x_{j}}\frac{\partial u_{i}}{\partial x_{j}}-\omega^{2}\right)$ (20)
@@
where the definition $S_{1}+S_{2}=1$ has been entered. The implication of Equation (21) is that if $\epsilon$ were known from the resolved scale, a relation between $S^{*}$ and $S_{1}$ is established for the rotational flamelet. This relation could be useful for the Flamelet Progress Variable model as well. For example, with $S_{1}=1/2$, we find $S^{*}=\sqrt{\epsilon/(3\nu)}$, its peak value. As $S_{1}$ varies in the range $0\leq S_{1}\leq 1$, $S^{*}$ varies narrowly between $\sqrt{\epsilon/(4\nu)}$ and $\sqrt{\epsilon/(3\nu)}$. In the range $1\leq S_{1}\leq 0$ (with inward swirl), $S^{*}$ varies more strongly between $\sqrt{\epsilon/(12\nu)}$ and $\sqrt{\epsilon/(3\nu)}$.
Viscous dissipation actually occurs over a range of the smallest scales and not totally at the Kolmogorov scale. Therefore, we have reason to multiply $\epsilon$ in Equations (20) and (21) by the coefficient $C_{vd}$ where $C_{vd}<1$. This yields
$C_{vd}\epsilon=4\nu S^{*2}[S_{1}^{2}+1-S_{1}]$ (22)
@@ 5 Kinetic Energy
With use of Equation (2), it can be readily found for the incoming streams of the flamelet that
@@
For the constant-density inflow portions of the counterflow structure, we can write for the inflow
@@
The right side of Equation (24) requires knowledge of three quantities for rotational flamelet theory: $S^{*},S_{1},$ and vorticity $\omega$. The Kolmogorov scale can be used to estimate the lengths. The kinetic energy per unit mass here is based on Kolmogorov scale velocity fluctuations; so, the value will be less than the statistical value for turbulence kinetic energy $k$. If a relation between the two kinetic energies is provided, knowledge of $k$ provides a relationship between those three quantities.
@@
We will use the relations from Pope, Page 185 [Pope2000]; the Kolmogorov velocity $u_{\kappa}\approx(\epsilon\nu)^{1/4}$ and the Kolmogorov length $\kappa\approx(\nu^{3}/\epsilon)^{1/4}$. We will average kinetic energy over a volume equal to $\kappa^{3}$. The last term in Equation (24) shows antisymmetry, taking positive values in two quadrants and negative values of identical magnitudes in the other two quadrants of any $\xi,\chi$ plane. Thereby, they have no consequence for our averaging. Equation (24) now yields after averaging the right side and relating the left side to $\epsilon$,
@@
$C_{ke}$ is a nondimensional coefficient of $O(1)$ that appears in the averaging process for the kinetic energy term on the left side of Equations (24). It accounts for both approximations made in averaging and adjusted decrease in kinetic energy at the Kolmogorov scale because dissipation has occurred over a range of the smallest scales. Realize that because the small-scale time scales are so short compared to filtered scale times, the associated averaging does not reduce final resolution. Thereby, Equations (21) and (25) have potential to be used with knowledge of $\epsilon$ from resolved-scale computation to determine the needed inflow parameters $S^{*}$ and $\omega$ for the rotational flamelet model.